Mary took 5 English tests. Her scores were 78%, 89%, 98%, 67%, and 90%.
What must she score on her sixth test so her average will be 85%?
: Mary took 5 English tests. Her scores were 78%, 89%, 98%, 67%, and 90%.
What must she score on her sixth test so her average will be 85%?

Answer by nerdybill(1120)

Let x = score on her sixth test
then to average an 85:
(78+89+98+67+90+x)/6 =85
(422+x)/6 =85
422+x = 510
x = 510 - 422
x = 88